OpenStudy (anonymous):

how do you do problems in math about belts and wheels such as the following: Two wheels are connected by a belt. The smaller wheel makes 20 revolutions every 3 seconds. The belt is 30 ft long and makes one complete loop every 8 seconds. Find the diameter of the smaller wheel.

OpenStudy (amistre64):

belts are linear and wheels are angular

OpenStudy (amistre64):

8*3 = 24 that should be a common time period to equate the 2

OpenStudy (amistre64):

20:3 as 30:8 *8 *3 ---------------- 160:24 as 90:24

OpenStudy (amistre64):

160 revolutions = 90 feet

OpenStudy (amistre64):

divide both sides by 160 160 revolutions = 90 feet /160 /160 ---------------------- 1 revolution = 9/16 feet

OpenStudy (amistre64):

youve got the circumference with that, so pull off the diamer using: C = pi d

OpenStudy (anonymous):

9/16pi ?

OpenStudy (amistre64):

with any luck .. yes :)

OpenStudy (anonymous):

thanks so much!

OpenStudy (amistre64):

hope its right ... it reads up right to me

ganeshie8 (ganeshie8):

i got the same answer. small wheel : 20 revs / 3 sec => 160 revs / 24 sec belt : 1 rev / 8 sec => 3 rev / 24 sec (or) 3*30 ft / 24 sec 160 revs/ 24 sec = 90 ft/ 24 sec 160 revs = 90 ft 1 rev = 9/16 ft also you may notice something interesting about BIG wheel. you can put any wheel that fits the loop. since it must be bigger than the smaller wheel, its diameter must be of range (9/16pi , 30-x)